OpenStudy (anonymous):

HEEELLLPPP! The speed of sound, V, near Earth's surface can be represented by this equation,v = 20√t + 273 where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number. Anwsers: 88, 90, 92, 94

OpenStudy (mathteacher1729):

This means: "solve the equation \(380 = 20\sqrt{t} +237\) "

OpenStudy (mathteacher1729):

What is the first step to solving that equation?

OpenStudy (anonymous):

IDK,.

OpenStudy (anonymous):

try, there is got to be a reason you are taking that class. if you do not know guess?

OpenStudy (anonymous):

basically you need to solve for t, which means make t on one side by itself.. so how would you go about that?

OpenStudy (anonymous):

Idk, i ask because ive rtied everything and CANT figure it out. I dont come on here to get critized i come on here for help.

OpenStudy (mathteacher1729):

@SKforlife when it comes to solving equations, is it the square root that's throwing you off in this example?

OpenStudy (anonymous):

I am trying to help you out, not criticizing you. But i can just solve the question for you and you won't learn a thing. If that is what you want then: \[380=20\sqrt{t}+237\] \[380-237=20\sqrt{t}\] \[143=20\sqrt{t}\] \[143/20=\sqrt{t}\] \[7.15=\sqrt{t}\] ; now square both sides. therefore t= 51.1225 the temperature when V = 380 meters per second to the nearest whole number is t= 51 degrees Celsius goodluck